# D ECOR : Improving Coherence in L2 English Writing with a Novel Benchmark for Incoherence Detection, Reasoning, and Rewriting

### v1

@@ Appendix C Additional statistics of DECOR
Appendix D Synthesizing training data with GPT-4
D.1 Dataset Synthesis
@@ D.2 Post-Processing
Appendix E Details of Experiments
E.1 Classification-based models
@@
E.2 Generation-based models
@@ E.3 Additional results
Models Context Current Sentence Reasons Condition Rewrite Llama2-7B So many investors struggled to make extraordinary shops. And they also have to care about the location of the shop and the menu. Because there are rival shops. But if someone tries a new thing, he doesn’t have to worry about the rival. Discourse marker w/ reason However, if someone tries a new thing, he doesn’t have to worry about the rival. w/o reason For example, if someone tries a new thing, he doesn’t have to worry about the rival. Llama3-8B -Instruct I found in several books of scientists, universities magazines, that people who want to succeed need to take risks, risk in research, risk in budgets, contracts, borrow investment. You can never do big money with the regular risks Semantic connection w/ reason You can never achieve big success without taking risks. w/o reason For example, you can never do big business with regular risks.
@@
Appendix F GPT-4 Prompts for Detection, Reasoning, and Rewriting
@@ F.5 Reasoning - Others
Appendix G GPT-4 Judge Prompt

### v2

@@ B.1 Inter-annotator agreement scores across different reason types
B.2 The achievability of high inter-annotator agreement on DECOR
We achieve a high inter-rater agreement for our dataset through our meticulously structured and clearly defined annotation process and scheme. Specifically, we recruited two expert annotators who are both professors with extensive experience in teaching English as a foreign language and have advanced degrees in Applied Linguistics. Before starting, we delivered a comprehensive tutorial and thoroughly reviewed our detailed annotation guidelines with the annotators. To familiarize them with our annotation scheme, we selected eight medium-level essays from TOEFL-11, generating 109 context-sentence pairs. The tutorial was structured into three sessions to ensure thorough coverage of all instances. During these sessions, annotators collaboratively labeled the samples. We also refined the annotation scheme as necessary. Additionally, we incorporated examples of errors encountered during these tutorial sessions into the annotation scheme. Subsequently, to quantitatively assess the inter-annotator agreement, we sampled an additional five medium-level essays, distinct from those used in the tutorial sessions, resulting in 72 context-sentence pairs.
Among the 72 samples, the first annotator labeled 35 context-sentence pairs as incoherent, and 30 as coherent. The second annotator labeled 39 pairs as incoherent, and 28 as coherent. Note that the rest were labeled as “uncertain”. Therefore, the resulting annotations for the 72 samples were relatively balanced. Given the high inter-annotator agreement achieved by the two annotators, with Cohen’s Kappa scores of 0.83 for incoherence detection and 0.90 for incoherence reasoning, we subsequently assigned them to independently annotate the entire test set.
@@ Appendix C Additional statistics of DECOR
Appendix D Human evaluation details
During human evaluation, our human expert was asked to compare system-generated rewrites with those produced by GPT-4. Specifically, we conducted three sets of pair-wise comparisons: 1) human-generated rewrites VS GPT-4; 2) rewrites generated by Llama2-7B trained with reasons VS GPT-4; and 3) rewrites generated by Llama2-7B trained without reasons VS GPT-4. For each set of pair-wise comparisons, for each pair of outputs (e.g. system 1 output VS system 2 output) the human expert was asked “Between system 1 and system 2 outputs, select the output that provides better rewrite based on the reason for incoherence; otherwise, choose “Tie” to indicate equal quality”.
After human evaluation, we interviewed the human expert and asked for his feedback on the system outputs. In general, the evaluator was surprised that our fine-tuned system can generate rewrites of decent quality. Among the three sets of comparisons, he was not able to tell which one was generated by humans. This indicates that our fine-tuned smaller language models are already capable of correcting the incoherence well based on the reasons behind it.
Appendix E Synthesizing training data with GPT-4
E.1 Dataset Synthesis
@@ E.2 Post-Processing
Appendix F Details of Experiments
F.1 Classification-based models
@@
F.2 Generation-based models
@@ F.3 Additional results
Models Context Current Sentence Reasons Condition Rewrite Llama2-7B So many investors struggled to make extraordinary shops. And they also have to care about the location of the shop and the menu. Because there are rival shops. But if someone tries a new thing, he doesn’t have to worry about the rival. Discourse marker w/ reason However, if someone tries a new thing, he doesn’t have to worry about the rival. w/o reason For example, if someone tries a new thing, he doesn’t have to worry about the rival. Llama3-8B -Instruct I found in several books of scientists, universities magazines, that people who want to succeed need to take risks, risk in research, risk in budgets, contracts, borrow investment. You can never do big money with the regular risks. Semantic connection w/ reason You can never achieve big success without taking risks. w/o reason For example, you can never do big business with regular risks.
@@
Appendix G GPT-4 Prompts for Detection, Reasoning, and Rewriting
@@ G.5 Reasoning - Others
Appendix H GPT-4 Judge Prompt